PROBLEM SET 5B 1. What are the formulas to calculate the probability of a un animous initial vote of Guilty or not-Guilty for the LA jurors in the criminal trial? Assume the probability of voting not-guity is 2/3 for each of the LA jurors. Use these formulas is explain why a very quick verdict in the OJ c riminal trial is very likely to be not- 2. Guilty
Answer #1

n = 12

P(not guilty), p = 2/3

P(guilty), q = 1 -2/3 = 1/3

P(X) = nCx px qn-x

1. Consider that there are 12 members in a Jury

P(unanimous verdict of Guilty) =(1/3)12

= 1.88x10-6

P(unanimous verdict of not guilty) = (2/3)12

= 0.0077

2. P(a very quick verdict of not guilty) = P(at least 7 Jury members say not guilty)

= 12C7x(2/3)7 x (1/3)5 + 12C8x(2/3)8 x (1/3)4 + 12C9 x (2/3)9 x (1/3)3 + 12C10 x (2/3)10 x (1/3)2 + 12x(2/3)11 x (1/3) + 0.0077

= 0.1908 + 0.2384 + 0.2120 + 0.1272 + 0.0462 + 0.0077

= 0.8223

So, there is high probability that quick verdict will be not guilty